find slope & y intercept for 2x-y+5=0 & 3x+4y=7

Answer by Alan3354(69443)

Solve for y. Then the slope is the coefficient of x.
The y-int is the constant.
---------
3x+4y = 7
4y = -3x + 7
y = (-3/4)x + 7/4
Slope m = -3/4
y-int = 7/4
------------
Do the other one the same way.